OpenStudy (anonymous):

Simplify the expression. (cos^2x + sin^2x)/(cot^2x - csc^2x)

OpenStudy (anonymous):

can I get a fan and medal if I get it correct?

OpenStudy (mathstudent55):

Hint: Pythagorean identities.

OpenStudy (princeharryyy):

the ans is 1

OpenStudy (princeharryyy):

they both are identities

OpenStudy (anonymous):

(cos^2x+sin^2x) / (csc^2x-cot^2x )= -(1/1)= -1

OpenStudy (cwrw238):

-1 is correct

OpenStudy (anonymous):

Thank you.

OpenStudy (cwrw238):

the numerator = 1 for the denominator use the identittyy csc^x = 1 + cot^2x

OpenStudy (princeharryyy):

yup -1 is correct i thought it was cosec^2x - cot^2x

OpenStudy (mathstudent55):

\(\color{red}{\sin^2 x + \cos^2 x = 1}\) Eq. 1 Divide both sides of Eq. 1 by \(\sin^2 x\): \(1 + \cot^2x = \csc^2 x\) \(\color{green}{\cot^2 x - \csc^2 x = -1}\) Eq. 2 Eq. 1 and Eq. 2 above are the identities you need. \(\dfrac{\color{red}{\cos^2 x + \sin^2 x}}{\color{green}{\cot^2 x - \csc^2 x}} = \dfrac{\color{red}{1}}{\color{green}{-1}} = -1\)

OpenStudy (anonymous):

(cos^2x + sin^2x)/(cot^2x - csc^2x) Fundamental identity shows: cos^2x+sin^2x=1 Plug in for (cos^2x+sin^2x) =1/(cot^2x-csc^2x) Fundamental identity shows: 1+cot^2x=csc^2x -csc^2x -csc^2x 1+cot^2x-csc^2=0 -1 -1 cot^2x-csc^2x=-1 Plug in for (cot^2x-csc^2x) =(1/-1) =-1